# Mathematical Modeling and Machine Learning for Predicting Shade-Seeking Behavior in Cows Under Heat Stress

### v1

@@ 3. Results of the model
On the one hand, the Decision Tree model has a RMSE of $16.03$, which is the highest among the models compared, meaning it is less accurate in prediction. However in terms of Interpretability, the Decision Tree scores Very High (that means the model is easy to understand), as its structure is simple and intuitive, resembling a flowchart. In a similar way the Explainability of the Decision Tree is rated as High, meaning that the decision-making process of the model can be clearly explained, making it easier to trace the reasoning behind predictions.
@@
Second, the plot reveals a few RMSE values extending toward the extremes (below 5 and above 25). This indicates that, in certain cases, the model performs less accurately. These outliers may correspond to specific scenarios or data points where the model struggles to generalize, possibly due to variations in environmental conditions or unmodeled factors. Addressing these cases could involve incorporating additional features or refining the model to improve its generalizability.
@@
(b) Model predictions for July 20th,
2023 ($Q3$).
@@
4. Conclusions
In this work, we have analyzed the performance of three Machine Learning models—Decision Trees, Random Forests, and Neural Networks—to predict the number of cows seeking shade as a response to varying environmental conditions. Using data from a farm in Titaguas, Valencia, the research aimed to determine which model could best predict shade-seeking behavior in response to the Temperature-Humidity Index (THI), a key indicator of thermal stress in cattle, while also exploring the capabilities and limitations of these models for livestock management under heat stress conditions.
@@ 4. Conclusions
In the article, several variables influencing shade-seeking behavior have been identified: in particular the time of day, current THI, and cumulative THI throughout the day and night. These factors strongly correlated with the cows’ movement towards or away from shaded areas, underscoring their relevance in managing thermal stress.
@@
In short, the mathematical frameworks applied in this article underline the critical role of rigorous modeling in addressing environmental factors affecting livestock productivity. By integrating Decision Trees, Random Forests, and Neural Networks with theoretically grounded metrics, this work not only enhances predictive accuracy but also establishes a versatile foundation for analyzing similar patterns in broader environmental systems. Future work may explore advanced methods such as dynamic system modeling or mathematical analysis to further expand the applicability of the presented models.

### v2

@@ 3.1. Global model performance
On the one hand, the Decision Tree model has a RMSE of $16.03$, which is the highest among the models compared, meaning it is less accurate for prediction. However in terms of Interpretability, the Decision Tree scores Very High (that means the model is easy to understand), as its structure is simple and intuitive, resembling a flowchart. In a similar way the Explainability of the Decision Tree is rated as High, meaning that the decision-making process of the model can be clearly explained, making it easier to trace the reasoning behind predictions.
@@ 3.2. Error distribution analysis
Second, the plot reveals a few RMSE values extending toward the extremes (and above 25). This indicates that, in certain cases, the model performs less accurately. These outliers may correspond to specific scenarios or data points where the model struggles to generalize, possibly due to variations in environmental conditions or unmodeled factors. Addressing these cases could involve incorporating additional features or refining the model to improve its generalizability.
3.3. Case studies
@@
(b) Model predictions for July 20th, 2023 ($Q3$).
@@
4. Discussion
In this work, we have analyzed the performance of three soft computing Machine Learning models—Decision Trees, Random Forests, and Neural Networks—to predict the number of cows seeking shade as a response to varying environmental conditions. Using data from a farm in Titaguas, Valencia, the research aimed to determine which model could best predict shade-seeking behavior in response to the Temperature-Humidity Index (THI), a key indicator of thermal stress in cattle, while also exploring the capabilities and limitations of these models for livestock management under heat stress conditions.
@@ 4. Discussion
The performance gap between the Random Forest model and the Neural Network is small (around $0.2$ RMSE), yet the tree-based ensemble is far easier to interpret and inherently robust to noisy or missing inputs—an essential property in commercial farms where cameras may be hidden and low-cost sensors drift over time—. This robustness makes the model an ideal upstream component for a hybrid soft-computing controller: its predicted shade-seeking count can feed a fuzzy-logic rule base that activates fans, sprinklers or retractable awnings according to linguistic rules such as: ”if THI is high and Predicted Shade is many, then increase airflow”. These neuro-fuzzy or RF-fuzzy combinations retain the data-driven accuracy of machine learning, while adding transparent, expert-defined control actions, providing a practical path to farm climate management.
In this article, several variables influencing shade-seeking behavior have been identified: in particular the time of day, current THI, and cumulative THI throughout the day and night. These factors strongly correlated with the cows’ movement towards or away from shaded areas, underscoring their relevance in managing thermal stress.
@@
5. Conclusions
This study demonstrates the potential of using soft computing approaches for mathematical modeling of noisy and highly variable biological behaviors. Using only climatic measurements and camera counts, both Random Forests and Neural Networks accurately predicted the number of dairy cows seeking shade during Mediterranean summer heat waves. The main conclusions are as follows:
Early warning capability::
the models anticipate shade-seeking peaks within one hour, with a median daily RMSE of $13.84$ cows.
Interpretability::
a $10$-tree Random Forest (depth $=5$) achieves an average RMSE of $14.9$ while retaining a transparent rule structure, making it the recommended choice for on-farm deployment.
Minimal feature set::
three easily derived thermal features—current THI, accumulated daytime THI and mean night-time THI—are sufficient for a low-cost decision-support system that can trigger ventilation, sprinkling or shading strategies in real time.
These results show that soft computing models provide robust, affordable tools for precision-livestock management aimed at mitigating heat stress and safeguarding animal welfare.